ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 47 Sharma, B. K. : Pt. Ravishankar Shukla University, Raipur, India. A Strong Convergence Theorem for Multivalued Nonexpansive Mappings A strong convergence result for approximating fixed points is proved for hybrid pair of multivalued nonexpansive mappings in reflexive Banach space with Gateaux differentiable norm [4]. Our result improves several known results [1], [2], [3] and [5].

[1] Browder, F. E. : Convergence of nonexpansives to fixed points of nonexpansive mappings in Banach spaces, Arch. Ration. Mach. Anal. 20 (1967)

[2] Helpern, B. : Fixed points of nonexpansive maps, Dull. Amer. Math. Soc. 73 (1967), 957-961.

[3] Reich, S. : Strong convergence theorems for resolvents of accretive operators in Banach spaces, J. Math. Anal. Appl. 75 (1980).

[4] Day, M. M. : Normed linear spaces,3rd Springer Verlag, Berlin/New York (1973).

[5] Latif, A. ,Husan, T. and Beg, T. : Fixed point of nonexpansive and k-multival ued maps, Internat. J. Math. Sci. 17(3) (1994), 429-436.